Loblaws has an inventory turnover rate of 12.3, an accounts payable period of 30 days, and...

  1. Loblaws has an inventory turnover rate of 12.3, an accounts payable period of 30 days, and an accounts receivable period of 40 days. What is the length of the cash cycle?

    39.67 days

    41.13 days

    37.88 days

    34.42 days

    30.71 days
Answer #1

Days in inventory=365/inventory turnover

=365/12.3=29.67 days

Cash cycle=Days in inventory+accounts receivable period-accounts payable period

=29.67+40-30 days

which is equal to

=39.67 days.